A toy cannon is anchored to the ceiling. It shoots a ball staight downward at -1.4 m/s . After 0.3 s , what is the displacement of the ball?
Answer #1

From Kinematic equation

Y=Yo + Vot - (1/2)gt2

when ball hits ground ,Y=0

0 =Yo+ (-1.4)(0.3) -(1/2)(9.8)(0.3)2

Yo=(1.4)(0.3) +(1/2)(9.8)(0.3)2

Yo = 0.861 m